OpenStudy (anonymous):

Suppose you know the graph of f contains the point (2,3). If g(x)=f(x+5). What point must be on the graph of g? Possible answers: (2,8) (-3,3) (7,2) (3,2) (2,-2)

OpenStudy (anonymous):

if \(g(x)=f(x+5)\) then \(g(-3)=f(-3+5)=f(2)=3\)

OpenStudy (anonymous):

So is the answer (-3,3)

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Ok so can you explain why f(2) = 3 in the end. That's the only part that confuses me